# Xiang Fan1,2    Sharath Girish2    Vivek Ramanujan1    Chaoyang Wang2    Ashkan Mirzaei2    Petr Sushko1    Aliaksandr Siarohin2    Sergey Tulyakov2    Ranjay Krishna1

### v1

@@ 4.2.1 Multi-view Static and Dynamic NVS
Table 3: Quantitative comparison on LLFF dataset. We outperform Stable Virtual Camera [zhou2025stable] significantly in terms of reconstruction quality for the static NVS task.
@@
Figure 4: Qualitative visualizations on a scene in the N3DV dataset. We consistently outperform [bai2025recammastercameracontrolledgenerativerendering, yu2025trajectorycrafter] on single view input while improving reconstruction quality with increasing number of input conditions.
We now extend our approach to the multi-view input setting consisting of images or time-synchronized videos. We first evaluate OmniView on the LLFF dataset [mildenhall2019locallff] consisting of multiple view captures of a static scene. We sparsely sample 3, 6, or 9 input views and choose one test view. We compare against SEVA [zhou2025stable] and GEN3C as they allow for multiple view inputs. Results are summarized in Tab. 3, measuring the 3 reconstruction metrics. We see that we perform competitively against the baselines across 3 metrics. While GEN3C is limited by its cache size which is 4 by default, we directly generalize to more input views despite being trained on upto 3 static views, while improving reconstruction quality with additional 3D information of the scene.
@@
Figure 5: Increasing number of conditional views improves reconstruction on N3DV [li2022neural] and is more aligned with the GT.
@@ 4.3 T2V/I2V + Camera Control
Table 4: Quantitative comparison on RE10K dataset. We continue to outperform [yu2025trajectorycrafter, ren2025gen3c] on the Image-to-Video + Camera Control task across all reconstruction metrics.
@@
We see that we obtain better reconstructions aligned with the GT views compared to [yu2025trajectorycrafter, ren2025gen3c] while also more closely aligning with the camera trajectory compared to AC3D [ac3d]. This is despite AC3D being trained primarily on the Camera Control task while we train on a variety of different 3D/4D tasks. Qualitative results are provided in the supplementary to show high video fidelity obtained even in the case of no input conditioning views.
@@ 6 Background: RoPE
RoPE [su2024roformer] injects positional information into self-attention by rotating each 2D channel pair of the query and key vectors with sinusoidal, index-dependent phases. This rotation makes attention scores a function of relative displacement (via phase differences) rather than absolute indices, yielding a smooth decay with distance, translation equivariance of the dot-product, and strong length/extrapolation behavior—all without adding parameters. For video tokens, the position is naturally 3D $(x,y,t)$; phases can be assigned per axis and combined (e.g., additively) to induce a spatio–temporal inductive bias. We use the shorthand $\text{RoPE}(\mathbf{q}_{xyt}^{z})=R(\mathbf{q}_{xyt}^{z},\boldsymbol{\theta}_{xyt})$ and apply the same operator to keys, where $\boldsymbol{\theta}_{xyt}$ denotes the sinusoidal phase parameters determined by $(x,y,t)$.
@@ 7 Ablation with PRoPE
In addition to the ablations shown with different variants in Table 6 of the main paper, we compare with another variation of RoPE for camera control [li2025camerasrelativepositionalencoding] (PRoPE). PRoPE consists of replacing the RoPE rotation matrix with the camera projection matrices for half of the dimensions while reorganizing the other half for 2D xy coordinates. They, however, require training the diffusion model from initialization as opposed to our approach which utilizes existing trained video diffusion models. Nevertheless, we evalute their approach in a finetuning setting using the pretrained video diffusion model. As they do not directly target the dynamic video setting, we extend their approach to modify the RoPE dimensions corresponding to the spatial coordinates while keeping the temporal dimensions as is. We train the model in this setting for 30K iterations and evaluate on the N3DV monocular video NVS task. Results are summarized in Tab. 7. We see that PRoPE leads to noisy reconstructions across all metrics likely due to the model incapable of significantly altering the underlying RoPE mechanism in the DiT, and can potentially require a large number of iterations for convergence. This is in contrast to our approach, which closely retains the 3D spatio-temporal RoPE structure, especially in the standard base setting of same cameras between different latents.
Figure 6: Qualitative visualizations on the static multi-view NVS task on the LLFF dataset Compared to Stable Virtual Camera [zhou2025stable], we obtain much higher detail along with aligned generations to the Ground-Truth, highlighted by the larger gap in SSIM values in Table 3 of the main paper.
@@
Figure 8: Qualitative visualizations on the I2V + Camera Control task on RE10K. Compared to prior works of [yu2025trajectorycrafter, ren2025gen3c], we obtain generations which are more 3D consistent with the input image condition and hence are better aligned with the Ground-Truth (Highlighted red boxes for example).
Figure 9: Qualitative visualizations on the T2V + Camera Control task on RE10K. Compared to AC3D [ac3d], we obtain generations which follow the camera trajectory in the source video while also producing high-fidelity video generations.
@@ 8 Qualitative visualizations
We visualize the generations from our approach as well as Stable Virtual Camera (SEVA) [zhou2025stable] for the case of 3 conditioning views and compare it with the Ground-Truth (GT) across 5 scenes in LLFF. Results are shown in Fig. 6. We see that we consistently obtain higher quality generations while SEVA leads to blurrier outputs. We also obtain generations which are more aligned with the GT such as the tip of the fern (top row), or the flower (second from top row), or the legs of the dinosaur (bottom row). We also visualize the generations with increasing number of views in Fig. 7. With increasing multi-views as input, the model reduces scale ambiguity reconstructing the scene more faithfully while also incorporating appearance information from additional input views.
Table 7: Ablation study comparing with PRoPE [li2025camerasrelativepositionalencoding] on N3DV.
@@
Next, we compare the camera control performance for the T2V setting with AC3D [ac3d]. Results are shown in Fig. 9. We utilize the camera trajectory from a source video and generate new videos with this trajectory. We see that AC3D can signficantly deviate from the source trajectory, such as the zooming in for the first scene while AC3D rotates left. We on the hand, are consistent with the camera input equivalently zooming into the scene. A similar but inverse result is observed in the third scene where the input camera slightly pans/rotates left while AC3D produces a zoomed in version, while we produce generations consistent with this trajectory.

### v2

@@ 4.2.1 Multi-view Static and Dynamic NVS
Table 3: Quantitative comparison on LLFF dataset. We outperform Stable Virtual Camera [91] significantly in terms of reconstruction quality for the static NVS task.
@@
Figure 4: Qualitative visualizations on a scene in the N3DV dataset. We consistently outperform [4, 83] on single view input while improving reconstruction quality with increasing number of input conditions.
We now extend our approach to the multi-view input setting consisting of images or time-synchronized videos. We first evaluate OmniView on the LLFF dataset [45] consisting of multiple view captures of a static scene. We sparsely sample 3, 6, or 9 input views and choose one test view. We compare against SEVA [91] and GEN3C as they allow for multiple view inputs. Results are summarized in Tab. 3, measuring the 3 reconstruction metrics. We see that we perform competitively against the baselines across 3 metrics. While GEN3C is limited by its cache size which is 4 by default, we directly generalize to more input views despite being trained on upto 3 static views, while improving reconstruction quality with additional 3D information of the scene.
@@
Figure 5: Increasing number of conditional views improves reconstruction on N3DV [33] and is more aligned with the GT.
@@ 4.3 T2V/I2V + Camera Control
Table 4: Quantitative comparison on RE10K dataset. We continue to outperform [83, 51] on the Image-to-Video + Camera Control task across all reconstruction metrics.
@@
We see that we obtain better reconstructions aligned with the GT views compared to [83, 51] while also more closely aligning with the camera trajectory compared to AC3D [2]. This is despite AC3D being trained primarily on the Camera Control task while we train on a variety of different 3D/4D tasks. Qualitative results are provided in the supplementary to show high video fidelity obtained even in the case of no input conditioning views.
@@ 6 Background: RoPE
RoPE [53] injects positional information into self-attention by rotating each 2D channel pair of the query and key vectors with sinusoidal, index-dependent phases. This rotation makes attention scores a function of relative displacement (via phase differences) rather than absolute indices, yielding a smooth decay with distance, translation equivariance of the dot-product, and strong length/extrapolation behavior—all without adding parameters. For video tokens, the position is naturally 3D $(x,y,t)$; phases can be assigned per axis and combined (e.g., additively) to induce a spatio–temporal inductive bias. We use the shorthand $\text{RoPE}(\mathbf{q}_{xyt}^{z})=R(\mathbf{q}_{xyt}^{z},\boldsymbol{\theta}_{xyt})$ and apply the same operator to keys, where $\boldsymbol{\theta}_{xyt}$ denotes the sinusoidal phase parameters determined by $(x,y,t)$.
@@ 7 Ablation with PRoPE
In addition to the ablations shown with different variants in Table 6 of the main paper, we compare with another variation of RoPE for camera control [32] (PRoPE). PRoPE consists of replacing the RoPE rotation matrix with the camera projection matrices for half of the dimensions while reorganizing the other half for 2D xy coordinates. They, however, require training the diffusion model from initialization as opposed to our approach which utilizes existing trained video diffusion models. Nevertheless, we evalute their approach in a finetuning setting using the pretrained video diffusion model. As they do not directly target the dynamic video setting, we extend their approach to modify the RoPE dimensions corresponding to the spatial coordinates while keeping the temporal dimensions as is. We train the model in this setting for 30K iterations and evaluate on the N3DV monocular video NVS task. Results are summarized in Tab. 7. We see that PRoPE leads to noisy reconstructions across all metrics likely due to the model incapable of significantly altering the underlying RoPE mechanism in the DiT, and can potentially require a large number of iterations for convergence. This is in contrast to our approach, which closely retains the 3D spatio-temporal RoPE structure, especially in the standard base setting of same cameras between different latents.
Figure 6: Qualitative visualizations on the static multi-view NVS task on the LLFF dataset Compared to Stable Virtual Camera [91], we obtain much higher detail along with aligned generations to the Ground-Truth, highlighted by the larger gap in SSIM values in Table 3 of the main paper.
@@
Figure 8: Qualitative visualizations on the I2V + Camera Control task on RE10K. Compared to prior works of [83, 51], we obtain generations which are more 3D consistent with the input image condition and hence are better aligned with the Ground-Truth (Highlighted red boxes for example).
Figure 9: Qualitative visualizations on the T2V + Camera Control task on RE10K. Compared to AC3D [2], we obtain generations which follow the camera trajectory in the source video while also producing high-fidelity video generations.
@@ 8 Qualitative visualizations
We visualize the generations from our approach as well as Stable Virtual Camera (SEVA) [91] for the case of 3 conditioning views and compare it with the Ground-Truth (GT) across 5 scenes in LLFF. Results are shown in Fig. 6. We see that we consistently obtain higher quality generations while SEVA leads to blurrier outputs. We also obtain generations which are more aligned with the GT such as the tip of the fern (top row), or the flower (second from top row), or the legs of the dinosaur (bottom row). We also visualize the generations with increasing number of views in Fig. 7. With increasing multi-views as input, the model reduces scale ambiguity reconstructing the scene more faithfully while also incorporating appearance information from additional input views.
Table 7: Ablation study comparing with PRoPE [32] on N3DV.
@@
Next, we compare the camera control performance for the T2V setting with AC3D [2]. Results are shown in Fig. 9. We utilize the camera trajectory from a source video and generate new videos with this trajectory. We see that AC3D can signficantly deviate from the source trajectory, such as the zooming in for the first scene while AC3D rotates left. We on the hand, are consistent with the camera input equivalently zooming into the scene. A similar but inverse result is observed in the third scene where the input camera slightly pans/rotates left while AC3D produces a zoomed in version, while we produce generations consistent with this trajectory.